OpenStudy (anonymous):

second-order linear ordinary differential equation y''-y = -2e^(-t) particulair solution yp = A*e^(-t) yp' = -A*e^(-t) yp'' = A*e^(-t) If i substitute it in the equation y''-y I get: A*e^(-t) - A*e^(-t) = -2*e^(-t) but the left side is zero?

OpenStudy (nikvist):

\[y_P=Ate^{-t}\]

OpenStudy (anonymous):

yeah but why?

OpenStudy (anonymous):

wolfram says it.. but how can you calculate it

OpenStudy (nikvist):

\[\mbox{because term }e^{-t}\mbox{ already exist in homogenic solution: }y_H=C_1e^t+C_2e^{-t}\]

myininaya (myininaya):

\[r^2-1=0 => r=\pm 1\]

OpenStudy (anonymous):

yeah i already got that one :)

OpenStudy (anonymous):

y = c1e^x + c2e^-x

OpenStudy (anonymous):

but the particulair one... drives me crazy

OpenStudy (anonymous):

it should be multiplied by t to make the particular solution linearly independent

myininaya (myininaya):

ok cool did you try \[y_p=Ae^{-t}+Bte^{-t}\] \[y_p'=-Ae^{-t}+Be^{-t}-Bte^{-t}\] \[y''_p=Ae^{-t}-Be^{-t}+Bte^{-t}-Be^{-t}\] so \[y''-y=e^{-t}(A-2B)+Bte^{-t}-[Ae^{-t}+Bte^{-t}]\] \[=e^{-t}(A-2B-A)+te^{-t}(B-B)=e^{-t}(-2B)\] this side is suppose to be equal to -2e^{-t} so -2B=-2 => B=1

OpenStudy (anonymous):

but ... then yp = (At+b)*e^(-t)? or yp = t*e^(-t) and do i have to take the derivative of this to substitute it

myininaya (myininaya):

and we didn't even need to plug in that homogeneous part i don't know why i did because that parts gonna turn out 0

myininaya (myininaya):

the second one jessica

myininaya (myininaya):

because remember be^{-t} is part of your homogenous solution

myininaya (myininaya):

its gonna end up zeroing out

myininaya (myininaya):

do you know what i mean?

OpenStudy (anonymous):

okee... so yp will be (at+b)e^(-t)

OpenStudy (anonymous):

just like what you did :) but a and b are vice versa

myininaya (myininaya):

you can do that jessica but what i'm saying is you only need ate^{-t} since be^{-t} is part of your homogenous solution

OpenStudy (anonymous):

ahhh oke.... thank you all for you explenations :)

myininaya (myininaya):

have a good day jessica i think i will be leaving soon

OpenStudy (anonymous):

yeaahh thank you :) i really appreciate your explanations :) i also wish you a good day :)